# Large Model Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v1

@@ VI-D Value Ecosystem of LM Agents
The creation of interconnected value network of LM agents empowers LM agents to autonomously and transparently manage value exchanges (e.g., data, knowledge, resources, and digital currencies), which is crucial for fostering innovation, enhancing cooperation, and driving economic growth within LM agents ecosystem. Blockchain technology provides a tamper-proof ledger that records all transactions between LM agents, ensuring transparency and trust in the system [101]. Smart contracts, which are self-executing agreements coded onto the blockchain, allow LM agents to autonomously manage transactions, enforce agreements, and execute tasks without the need for intermediaries [231]. Additionally, the integration of oracles—trusted data sources that feed real-world information into the blockchain—enables LM agents to interact with external data and execute contracts based on real-time conditions, further enhancing the functionality of value networks.
@@ VII Conclusion
In this paper, we have provided an in-depth survey of the state-of-the-art in the architecture, interaction paradigms, security and privacy, and future trends of LM agents. Specifically, we have introduce a novel architecture and its key components, critical characteristics, enabling technologies, and potential applications, toward embodied, autonomous, and connected intelligence of LM agents. Afterward, we have explored the taxonomy of interaction patterns and practical collaboration paradigms among LM agents, including data, computation, and information sharing for collective intelligence. Furthermore, we have identified significant security and privacy threats inherent in the ecosystem of LM agents, discussed the challenges of security/privacy protections in multi-agent environments, and reviewed existing and potential countermeasures. As the field progresses, ongoing research and innovation will be crucial for overcoming existing limitations and harnessing the full potential of LM agents in transforming intelligent systems.

### v2

@@ VI-D Value Ecosystem of LM Agents
The creation of interconnected value network of LM agents empowers LM agents to autonomously and transparently manage value exchanges (e.g., data, knowledge, resources, and digital currencies), which is crucial for fostering innovation, enhancing cooperation, and driving economic growth within LM agents ecosystem. Blockchain technology provides a tamper-proof ledger that records all transactions between LM agents, ensuring transparency and trust in the system. Smart contracts, which are self-executing agreements coded onto the blockchain, allow LM agents to autonomously manage transactions, enforce agreements, and execute tasks without the need for intermediaries [207]. Additionally, the integration of oracles, i.e., trusted data sources that feed real-world information into the blockchain, enables LM agents to interact with external data and execute contracts based on real-time conditions, further enhancing the functionality of value networks.
@@ VII Conclusion
In this paper, we have provided an in-depth survey of the state-of-the-art in the architecture, cooperation paradigms, security and privacy, and future trends of LM agent networks. Specifically, we have introduced a novel architecture and its key components, critical characteristics, enabling technologies, and potential applications, toward autonomous, embodied, and connected intelligence of LM agents. Afterward, we have explored the taxonomy of interaction patterns and practical collaboration paradigms among LM agents, including data, computation, and information sharing for collective intelligence. Furthermore, we have identified significant security and privacy threats inherent in the ecosystem of LM agents, discussed the challenges of security/privacy protections in multi-agent environments, and reviewed existing and potential countermeasures. As the field progresses, ongoing research and innovation will be crucial for overcoming existing limitations and harnessing the full potential of LM agents in transforming intelligent systems.